Find the rejection region for a test of independence of two classifications where the contingency table contains r rows and c columns. A. r = 3, c = 6, alpha = 0.10
Answer #1

df = (r - 1) * (c - 1) = 2 * 5 = 10

Test statistic

Rejection region is test statistic > 15.987